OpenStudy (anonymous):

what is percent 224 of 56?

rebeccaxhawaii (rebeccaxhawaii):

what is 24% or 56?

rebeccaxhawaii (rebeccaxhawaii):

.24x56=?

OpenStudy (anonymous):

125.44

rebeccaxhawaii (rebeccaxhawaii):

please lead the OS user to understand the material rather than giving direct answers and you must explain BEFORE giving out the answer so they are able to use this knowledge long-term so please help lead them to the correct answer

jhonyy9 (jhonyy9):

so you think 224% of 56 ?

jhonyy9 (jhonyy9):

so than you can using the rule of 3 simple you know that 56 is 100% so than you can writing 56 ---------- 100% x ---------- 224% cross multiplie and calculi the value of x ------------------------- 100*x = 224*56 x = ?